OpenStudy (anonymous):

Paula spots a glider located at an angle of elevation of 42°. The distance between the glider and Paula is 3,280 feet. To the nearest foot, what is the height of the glider h from the ground? Show your work

OpenStudy (idkwut):

h / 3280 = sin(42) h = 3280(sin(42)) = ? Solve.

OpenStudy (anonymous):

2194.75?

OpenStudy (idkwut):

Good job! :D

OpenStudy (anonymous):

QvQ thank you for your help